How to graph a function of the for f(x)+f(X)on the same curvelike for eg : to find the minimum value of sinx+cosx

Grade:12th pass

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,

f(x) = sin x + cos x
sin (a + b) = sin a.cos b + cos a.sin b
sin x + cos x =2^1/2 sin (x +π/4)
f(x) =2^1/2 sin (x +π/4)
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty